# Static and Dynamic BART for Rank-Order Data

### v4

@@ 1 Introduction
Rank-order data emerge when an expert is asked to rank a finite set of $N$ items according to a certain criterion, such as personal preference or a predetermined statistic. The study of rank-order data is becoming popular in several fields, including political science (Gormley and Murphy, 2008), music preferences (Bradlow and Fader, 2001), economics (Fok et al., 2012), and sports analytics (Graves et al., 2003). In practice, the rankers are likely heterogeneous in their qualities or opinions, and it is often of interest to investigate how the rankings may be dependent on covariates of the ranked items (e.g., summary statistics). For example, for sports leagues such as the NCAA, NFL, or NBA, many experts provide ranking lists of teams by relying on their experiences and team-specific statistics, like the win percentage and the margin of victory. Moreover, reflecting the inherently dynamic nature of competitors’ abilities and human preferences, the experts’ rankings may change over time (e.g., Graves et al., 2003). Analogous considerations hold for rank-order data from other fields, such as marketing, politics, and economics.
In this article, we focus on models for rank-order data (Alvo and Philip, 2014; Liu et al., 2019). Specifically, we consider the family of Thurstone models, which assumes the existence of an unobserved evaluation score for each of the $N$ items whose noisy realizations determine the rankings provided by the $M$ experts. However, most of the existing models in this class are designed to study cross-sectional data and are inherently static, thus ignoring any temporal dependence.
@@
Figure 1 shows the time-varying rankings of the Economic Complexity Index (ECI) (Simoes and Hidalgo, 2011) for 8 countries from 1965 to 2015, and the 2022 weekly ranking of NCAA teams by a single pollster. The changes in rankings over time are heterogeneous in magnitude, with both large and small shifts, as well as a few series having a stable ranking over time. Overall, these features highlight both slow evolution and large breaks due to events such as global crises or small events at a country level. The NCAA dataset shows similar features, with apparently smoother dynamics due to the smaller number of items. Section 5 includes an application to these datasets among the others.
@@
This article contributes to the literature on rank-order models by proposing two new methods for addressing these issues. Specifically, we leverage the Bayesian additive regression trees (BART, see Chipman et al., 2010) framework to define a flexible nonparametric static model for rank-order data, named rank-order Bayesian additive regression trees (ROBART). This model is an extension of the linear rank-order model of Li et al. (2022) to a setting with non-linearities and covariates that are specific to both items and rankers. Then, we extend this framework to a dynamic model called autoregressive rank-order Bayesian additive regression trees (ARROBART), which accounts for the persistence in time series of rank-order data (e.g., see Fig. 1).
@@
The most widely used families of models for rank data are the Thurstone (Thurstone, 1927) and Plackett-Luce (Luce, 1959; Plackett, 1975) models, which differ in the distribution of the noise in the random utility representations (normal and Gumbel, respectively) and in the ordering interpretation (joint or sequential). Relying on normal regression models makes the Thurstone family easier to interpret and extend in several directions. However, the estimation of these models is generally difficult, especially for large numbers of ranked items, due to the high-dimensional integral form of the likelihood function. Within the Thurstone family, the Bayesian approach coupled with data augmentation methods provides an advantageous computational alternative that also allows for the introduction of covariates (Yu, 2000). Johnson et al. (2002) developed a hierarchical Bayesian Thurstone model for rank-order data from known groups of evaluators, whereas Li et al. (2022) introduced ranker-specific error variances and a Dirichlet process mixture prior for ranker-specific effects of covariates. Recently, Gu and Yu (2023) introduced a Bayesian spatial autoregressive model to learn ranking data in social networks and to account for social dependencies among individuals. We contribute to this literature by proposing two new Thurstone models, a static and a dynamic one, where a Bayesian nonparametric approach is used to capture the possibly nonlinear impact of covariates. Moreover, we extend the proposed methods to address incomplete data (i.e., partial rankings) in the form of top-$q$ rankings, which frequently happens when rankers assess many items.
@@
The study of time series of rankings is less developed but has attracted increasing attention in recent years. Most existing dynamic ranking models are generalizations of the Plackett-Luce model (Bradlow and Fader, 2001; Grewal et al., 2008; Glickman and Hennessy, 2015; Henderson and Kirrane, 2018), whereas this article provides a dynamic extension to the Thurstone family. We propose a novel time-varying nonparametric approach that considers an autoregressive process for the latent scores and uses a specification based on regression trees to flexibly model the impact of the lagged response and the external covariates. Its linear counterpart is also introduced as a benchmark.
@@
The possibility of having large deviations as well as small changes in the time series of ranking lists, as shown in Fig. 1, calls for the use of flexible tools to model the dynamics of the latent “scores” driving the ranking of each item. To address these issues, we introduce the BART specification to model the conditional mean of latent scores. We leverage the good performance of this approach for fitting and forecasting regression models (e.g., see Hill et al., 2020) in order-statistics models to design static and dynamic nonparametric frameworks for rank-order data. Despite its advantages, BART has rarely been used in time series or state space models. Recent advances include Bayesian additive vector autoregressive trees for multiple time series (Huber and Rossini, 2022), and an extension to a mixed-frequency context for nowcasting in the presence of extreme observations (Huber et al., 2023) . Our contribution to this literature concerns the derivation of new closed-form filtering, predictive, and smoothing distributions for the latent time-varying scores in the ARROBART model. Furthermore, we show the proportionality of these distributions to finite mixtures.
@@ 2 Nonparametric Thurstone Models for Rank-Order Data
This section introduces notation, then presents a new class of nonparametric models for rank-order data that extends the Thurstone model (Thurstone, 1927) by allowing for flexible, possibly nonlinear functions of the covariates. Then, we extend the framework to account for temporal dependence and persistence for investigating and forecasting time series data.
A full ranking of $N$ items is a map $\tau:\mathds{A}\to\mathcal{P}_{N}$ from a finite, ordered set of items, $\mathds{A}=\{a_{1},\ldots,a_{N}\}$ to the space of $N$-dimensional permutations, $\mathcal{P}_{N}$. Let ##\## denote the cardinality of a set. With a slight abuse of notation, we denote a full ranking of $N$ items by the vector $\tau=(\tau_{1},\ldots,\tau_{N})^{\prime}\in\mathcal{P}_{N}$, where $\tau_{i}=\tau(i)$ is the rank of item $a_{i}\in\mathds{A}$, for $i=1,\ldots,N$, and $\tau_{i}<\tau_{j}$ means that item $a_{i}$ is preferred to $a_{j}$, written $a_{i}\succ a_{j}$ since it is assigned a lower rank. Rankings and pairwise comparisons are related, as, given a full ranking $\tau$, it holds $[a_{i}\succ a_{j}]\iff\tau_{i}<\tau_{j}$ for each $i,j\in\{1,\ldots,N\}$. Therefore, a full ranking $\tau$ is an ordered $N$-tuple representing a set of non-contradictory pairwise relationships in $\mathds{A}$, and as such, it encodes an ordered preference for all the items in $\mathds{A}$.
In this article, we consider a setting where $N$ items are ordered by $M$ rankers. For each item $i=1,\ldots,N$ and ranker $j=1,\ldots,M$, the $K_{x}$-dimensional vector of covariates is denoted by $\mathbf{X}_{ij}\in\mathds{R}^{K_{x}}$. To account for external information that might affect the rankers’ decision process, it is possible to include item- and ranker-specific covariates, $\mathbf{x}_{i}^{a}\in\mathds{R}^{K_{a}}$ and $\mathbf{x}_{j}^{r}\in\mathds{R}^{K_{r}}$, as well as covariates specific to both items and rankers, $\mathbf{w}_{ij}\in\mathds{R}^{K_{w}}$. In the dynamic case, this setup is replicated for each of the $T$ time periods. Finally, for any vector $\mathbf{z}=(z_{1},\ldots,z_{N})^{\prime}\in\mathds{R}^{N}$, we use $\operatorname{rank}(\mathbf{z})=[z_{i_{1}}\succ\ldots\succ z_{i_{N}}]$ to denote the full ranking of $z_{i}$’s in decreasing order of preference, that is, $z_{i_{1}}<\ldots<z_{i_{N}}$.
@@ 2.2 Static Model: Rank-Order BART
The Thurstone model can be described as a random utility model (e.g., see Walker and Ben-Akiva, 2002), where the full ranking $\tau$ is driven by a latent score, $\mathbf{z}$, such that for any two items $i_{1},i_{2}$, it holds $z_{i_{1}}<z_{i_{2}}$ if and only if $i_{1}\succ i_{2}$. The latent vector $\mathbf{z}$ is assumed to follow a multivariate normal distribution with mean $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{N})^{\prime}$ representing the score of each item:
@@
[…]
Consider a framework where no temporal observations are available or the time dependence is irrelevant to the analysis. Moreover, we assume each individual in a collection of $M$ professionals or rankers provides a ranking list of the same set of $N$ items, as is common in many real-world situations such as sports analytics and customer reviews. To investigate data without temporal dependence, we define a new static model that extends the Thurston approach in Eq. (2) by including the BART specification from Eq. (1) for the conditional mean of the latent scores. Specifically, we generalise Eq. (2) by assuming ranker-specific scores, $\mathbf{z}_{j}$, and, for each pair $(i,j)$, we replace the intercept $\gamma_{ij}$ with an unknown function of the covariates, $f$. Moreover, to cope with potential nonlinearities, interactions, and discontinuities in the rankers’ dynamic scores, we adopt a nonparametric approach based on a BART specification for the latent regression function, $f$.
@@
[…]
@@ 2.3 Dynamic Model: AutoRegressive Rank-Order BART
where the regression function $f$ is approximated by a BART specification. A special case of the ARROBARTX in Eq. (5) is an AutoRegressive Rank-Order BART model without covariates (ARROBART) and with one lag, which is obtained by assuming $\mathbf{X}_{ij,t}=\mathbf{z}_{j,t-1}$. The ARROBART model can be interpreted as a hidden Markov model (HMM, see Cappé et al., 2005) with measurement and transition density given by, respectively:
@@
The temporal dependence among the latent variables introduces a filtering problem in the estimation of the ARROBART model, as making inferences on the latent vector $\mathbf{z}_{t}$ requires the derivation of appropriate filtering and smoothing distributions.
Let us first introduce some additional notation. For a variable $x\in\mathcal{D}\subseteq\mathds{R}$, the BART specification induces a partition $\{C_{k}\}_{k=1}^{K}$ of the domain $\mathcal{D}$, such that the (approximated) function has constant value $\tilde{\mu}_{k}$ over each region $C_{k}$. We refer to the Supplement for proof and give a sketch below. First, notice that, for each tree $s\in\{1,\ldots,S\}$ defined in eq. (1), the collection of intervals $\{B_{\ell,s}\}_{\ell=1}^{b_{s}}$ is a partition of $\mathcal{D}$ and the $\ell$th leaf of the $s$th tree is associated to the scalar coefficient $\mu_{\ell,s}\in\mathds{R}$. Moreover, we can associate each tuple (j1,,jS)
@@
$p(\mathbf{z}_{t}|\tau_{1:t})\propto\mathbb{I}(\mathbf{z}_{t}\in A_{t})\sum_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{k}\in\{1,\ldots,K\}^{N}}\mathcal{N}_{N}(\mathbf{z}_{t}|\bm{\tilde{\mu}}_{\mathbf{k}},I_{N})q_{\mathbf{k},t}(A_{1:t-1}),$ (7)
@@
$q_{\mathbf{k},t}(A_{1:t-1})$ $=\sum_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{m}\in\{1,\ldots,K\}^{N}}q_{\mathbf{m},t-1}(A_{1:t-2})\int_{\mathbf{C}_{\mathbf{k},t-1}}\mathcal{N}_{N}(\mathbf{z}_{t-1}|\bm{\tilde{\mu}}_{\mathbf{m}},I_{N})\>\mathrm{d}\mathbf{z}_{t-1},$
@@
$p(\mathbf{z}_{t+1}|\tau_{1:t})\propto\sum_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{k}\in\{1,\ldots,K\}^{N}}\mathcal{N}_{N}(\mathbf{z}_{t+1}|\bm{\tilde{\mu}}_{\mathbf{k}},I_{N})q_{\mathbf{k},t+1}(A_{1:t}).$ (8)
@@
Corollary 1 shows that both the filtering and smoothing distributions are proportional to finite mixtures. These alternative representations for the filtering and smoothing distributions can be used to design a more efficient algorithm for posterior inference based on a data augmentation approach coupled with a Metropolis-Hastings step (see Section 3).
@@
$p(\mathbf{z}_{t}|\tau_{1:t})$ $\propto\sum_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{k}\in\{1,\ldots,K\}^{N}}w^{F}_{\mathbf{k},t}u^{F}_{\mathbf{k},t}(\mathbf{z}_{t}),$ (9)
@@
$w^{F}_{\mathbf{k},t}$ $=\frac{q_{\mathbf{k},t}(A_{1:t-1})n_{\mathbf{k},t}}{\sum_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{m}\in\{1,\ldots,K\}^{N}}q_{\mathbf{m},t}(A_{1:t-1})n_{\mathbf{m},t}}.$
@@
$p(\mathbf{z}_{t}|\tau_{1:T})$ $\propto\sum_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{k}\in\{1,\ldots,K\}^{N}}w^{S}_{\mathbf{k},t}u^{S}_{\mathbf{k},t}(\mathbf{z}_{t}),$ (10)
@@
$w^{S}_{\mathbf{k},t}$ $=\frac{q_{\mathbf{k},t}(A_{1:t-1})m_{\mathbf{k},t}}{\sum_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\ell}\in\{1,\ldots,K\}^{N}}q_{\bm{\ell},t}(A_{1:t-1})m_{\bm{\ell},t}},\qquad\mathbf{k}\in\{1,\ldots,K\}^{N}.$

### v5

@@ 1 Introduction
Rank-order data emerge when an expert is asked to rank a finite set of $N$ items according to a certain criterion, such as personal preference or a predetermined statistic. The study of rank-order data is becoming popular in several fields, including political science (gormley2008exploring), music preferences (bradlow2001bayesian), economics (FokPaapROL2012), and sports analytics (graves2003hierarchical). In practice, the rankers are likely heterogeneous in their qualities or opinions, and it is often of interest to investigate how the rankings may be dependent on covariates of the ranked items (e.g., summary statistics). For example, for sports leagues such as the NCAA, NFL, or NBA, many experts provide ranking lists of teams by relying on their experiences and team-specific statistics, like the win percentage and the margin of victory. Moreover, reflecting the inherently dynamic nature of competitors’ abilities and human preferences, the experts’ rankings may change over time (e.g., graves2003hierarchical). Analogous considerations hold for rank-order data from other fields, such as marketing, politics, and economics.
In this article, we focus on models for rank-order data (alvo2014statistical; liu2019model). Specifically, we consider the family of Thurstone models, which assumes the existence of an unobserved evaluation score for each of the $N$ items whose noisy realizations determine the rankings provided by the $M$ experts. However, most of the existing models in this class are designed to study cross-sectional data and are inherently static, thus ignoring any temporal dependence.
@@
Figure 1 shows the time-varying rankings of the Economic Complexity Index (ECI) (simoes2011economic) for 8 countries from 1965 to 2015, and the 2022 weekly ranking of NCAA teams by a single pollster. The changes in rankings over time are heterogeneous in magnitude, with both large and small shifts, as well as a few series having a stable ranking over time. Overall, these features highlight both slow evolution and large breaks due to events such as global crises or small events at a country level. The NCAA dataset shows similar features, with smoother dynamics. Section 5 includes an application to these datasets among the others.
@@
This article contributes to the literature on rank-order models by proposing two new methods for addressing these issues. Specifically, we leverage the Bayesian additive regression trees (BART, see chipman2010bart) framework to define a flexible nonparametric static model for rank-order data, named rank-order Bayesian additive regression trees (ROBART). This model is an extension of the linear rank-order model of li2016bayesian to a setting with non-linearities and covariates that are specific to both items and rankers. Then, we extend this framework to a dynamic model called autoregressive rank-order Bayesian additive regression trees (ARROBART), which accounts for the persistence in time series of rank-order data (e.g., see Fig. 1).
@@
The most widely used families of models for rank data are the Thurstone (thurstone1927law) and Plackett-Luce (luce1959individual; plackett1975analysis) models, which differ in the distribution of the noise in the random utility representations (normal and Gumbel, respectively) and in the ordering interpretation (joint or sequential). Relying on normal regression models makes the Thurstone family easier to interpret and extend in several directions. However, the estimation of these models is generally difficult, especially for large numbers of ranked items, due to the high-dimensional integral form of the likelihood function. Within the Thurstone family, the Bayesian approach coupled with data augmentation methods provides an advantageous computational alternative that also allows for the introduction of covariates (yu2000bayesian). johnson2002bayesian developed a hierarchical Bayesian Thurstone model for rank-order data from known groups of evaluators, whereas li2016bayesian introduced ranker-specific error variances and a Dirichlet process mixture prior for ranker-specific effects of covariates. Recently, gu2023social introduced a Bayesian spatial autoregressive model to learn ranking data in social networks and to account for social dependencies among individuals. We contribute to this literature by proposing two new Thurstone models, a static and a dynamic one, where a Bayesian nonparametric approach is used to capture the possibly nonlinear impact of covariates. Moreover, we extend the proposed methods to address incomplete data (i.e., partial rankings) in the form of top-$q$ rankings, which frequently happens when rankers assess many items.
@@
The study of time series of rankings is less developed but has attracted increasing attention in recent years. Most existing dynamic ranking models are generalizations of the Plackett-Luce model (bradlow2001bayesian; grewal2008university; glickman2015stochastic; henderson2018comparison), whereas this article provides a dynamic extension to the Thurstone family. We propose a novel time-varying nonparametric approach that considers an autoregressive process for the latent scores and uses a specification based on regression trees to flexibly model the impact of the lagged response and the external covariates. Its linear counterpart is also introduced as a benchmark.
@@
The possibility of having large deviations as well as small changes in the time series of ranking lists, as shown in Fig. 1, calls for the use of flexible tools to model the dynamics of the latent “scores” driving the ranking of each item. To address these issues, we introduce the BART specification to model the conditional mean of latent scores. We leverage the good performance of this approach for fitting and forecasting regression models (e.g., see hill2020bayesian) in order-statistics models to design static and dynamic nonparametric frameworks for rank-order data. Despite its advantages, BART has rarely been used in time series or state space models. Recent advances include Bayesian additive vector autoregressive trees for multiple time series (huber2021inference), and an extension to a mixed-frequency context for nowcasting in the presence of extreme observations (huber2023nowcasting) . Our contribution to this literature concerns the derivation of new closed-form filtering, predictive, backward sampling, and smoothing distributions for the latent time-varying scores in the ARROBART model. Furthermore, we show the proportionality of these distributions to finite mixtures.
@@ 2 Nonparametric Thurstone Models for Rank-Order Data
This section introduces notation, then presents a new class of nonparametric models for rank-order data that extends the Thurstone model (thurstone1927law) by allowing for flexible, possibly nonlinear functions of the covariates. Then, we extend the framework to account for temporal dependence and persistence for investigating and forecasting time series data.
A full ranking of $N$ items is a map $\tau:\mathds{A}\to\mathcal{P}_{N}$ from a finite, ordered set of items, $\mathds{A}=\{a_{1},\ldots,a_{N}\}$ to the space of $N$-dimensional permutations, $\mathcal{P}_{N}$. Let #\# denote the cardinality of a set. With a slight abuse of notation, we denote a full ranking of $N$ items by the vector $\tau=(\tau_{1},\ldots,\tau_{N})^{\prime}\in\mathcal{P}_{N}$, where $\tau_{i}=\tau(i)$ is the rank of item $a_{i}\in\mathds{A}$, for $i=1,\ldots,N$, and $\tau_{i}<\tau_{j}$ means that item $a_{i}$ is preferred to $a_{j}$, written $a_{i}\succ a_{j}$ since it is assigned a lower rank. Rankings and pairwise comparisons are related, as, given a full ranking $\tau$, it holds $a_{i}\succ a_{j}]\iff\tau_{i}<\tau_{j}$ for each $i,j\in\{1,\ldots,N\}$. Therefore, a full ranking $\tau$ is an ordered $N$-tuple representing a set of non-contradictory pairwise relationships in $\mathds{A}$, and as such, it encodes an ordered preference for all the items in $\mathds{A}$.
In this article, we consider a setting where $N$ items are ordered by $M$ rankers. For each item $i=1,\ldots,N$ and ranker $j=1,\ldots,M$, the $K_{x}$-dimensional vector of covariates is denoted by $\mathbf{X}_{ij}\in\mathds{R}^{K_{x}}$. To account for external information that might affect the rankers’ decision process, it is possible to include item- and ranker-specific covariates, $\mathbf{x}_{i}^{a}\in\mathds{R}^{K_{a}}$ and $\mathbf{x}_{j}^{r}\in\mathds{R}^{K_{r}}$, as well as covariates specific to both items and rankers, $\mathbf{w}_{ij}\in\mathds{R}^{K_{w}}$. In the dynamic case, this setup is replicated for each of the $T$ time periods. Finally, for any vector $\mathbf{z}=(z_{1},\ldots,z_{N})^{\prime}\in\mathds{R}^{N}$, we use $\operatorname{rank}(\mathbf{z})=[z_{i_{1}}\succ\ldots\succ z_{i_{N}}]$ to denote the full ranking of $z_{i}$’s in decreasing order of preference, that is, $z_{i_{1}}<\ldots<z_{i_{N}}$, i.e. lower latent variable scores correspond to higher rankings (more preferred or “better” items).
@@ 2.2 Static Model: Rank-Order BART
The Thurstone model can be described as a random utility model (e.g., see walker2002generalized), where the full ranking $\tau$ is driven by a latent score, $\mathbf{z}$, such that for any two items $i_{1},i_{2}$, it holds $z_{i_{1}}<z_{i_{2}}$ if and only if $i_{1}\succ i_{2}$. The latent vector $\mathbf{z}$ is assumed to follow a multivariate normal distribution with mean $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{N})^{\prime}$ representing the score of each item:
@@
The model in Eq. (S.5) poses an identification problem since the observations only provide information about the ordering of the elements in $\mathbf{z}$, which implies that the parameters $(\bm{\gamma},\sigma_{\varepsilon}^{2})$ are not separately identifiable from the likelihood. Therefore, to ensure the identification of the parameters, we fix $\sigma_{\varepsilon}^{2}=1$. [^1] If one is interested in interpreting the coefficients of some covariates, then it would be necessary to impose the additional constraint the $\bm{\gamma}\in\Gamma=\{\bm{\gamma}\in\mathds{R}^{N}:\mathbf{1}_{N}^{\prime}\bm{\gamma}=0\}$, where $\mathbf{1}_{N}$ is an $N$-dimensional vector with all entries equal to 1.
Consider a framework where no temporal observations are available or the time dependence is irrelevant to the analysis. Moreover, we assume each individual in a collection of $M$ professionals or rankers provides a ranking list of the same set of $N$ items, as is common in many real-world situations such as sports analytics and customer reviews. To investigate data without temporal dependence, we define a new static model that extends the Thurston approach in Eq. (S.5) by including the BART specification from Eq. (1) for the conditional mean of the latent scores. Specifically, we generalise Eq. (S.5) by assuming ranker-specific scores, $\mathbf{z}_{j}$, and, for each pair $(i,j)$, we replace the intercept $\gamma_{ij}$ with an unknown function of the covariates, $f$. Moreover, to cope with potential nonlinearities, interactions, and discontinuities in the rankers’ dynamic scores, we adopt a nonparametric approach based on a BART specification for the latent regression function, $f$.
@@
where $f:\mathds{R}^{K_{x}}\to\mathds{R}$ is a regression function for which we assume a BART specification. Denoting with $\delta_{\mathbf{z}}(\tau)$ the Dirac mass at $\mathbf{z}$, $\left\lVert\cdot\right\rVert$ the $L_{2}$ norm, and $\bm{\theta}$ a vector collecting all the static parameters (ranker- and item-specific), one obtains the likelihood:
@@ 2.3 Dynamic Model: AutoRegressive Rank-Order BART
where the regression function $f$ is approximated by a BART specification. A special case of the ARROBARTX in Eq. (5) is an AutoRegressive Rank-Order BART model without covariates (ARROBART) and with one lag, which is obtained by assuming $\mathbf{X}_{ij,t}=\mathbf{z}_{j,t-1}$. The ARROBART model can be interpreted as a hidden Markov model (HMM, see cappe2005hidden) with measurement and transition density given by, respectively:
@@
The temporal dependence among the latent variables introduces a filtering problem in the estimation of the ARROBART model, as making inferences on the latent vector $\mathbf{z}_{t}$ requires the derivation of appropriate filtering distributions and appropriate sampling strategies.
[…]
@@
For partial ranking data, if missingness of a rank implies absence of knowledge of an inequality between elements of $\bm{z}_{t}$, then the set $A_{t}$ is not constrained by the unknown inequality. In detail, partial ranking occurs when each assessor provides the (full) ranking of a subset $A\subset\mathbb{A}=\{a_{1},\ldots,a_{N}\}$ consisting of $k\leq N$ items, while leaving the other items $a_{\ell}\in A^{c}$ not ranked. Let us denote this ranking of subset $A$ by $\tilde{\bm{\tau}}$, then we define a partial ranking as $\bm{\tau}:\mathbb{A}\to\mathcal{S}\subset\mathcal{P}_{N}$, where $\mathcal{S}=\{\bm{\tau}\in\mathcal{P}_{N}:\text{ if }\tilde{\tau}_{i_{1}}<\tilde{\tau}_{i_{2}},\text{ with }a_{i_{1}},a_{i_{2}}\in A\text{ then }\tau_{i_{1}}<\tau_{i_{2}}\}$ is the set of possible augmented random permutation vectors, which coincides with the original partially ranked items together with the allowable values of the missing ranks. This means that the order between items $a_{i}\in A$ in the latent rank vector $\bm{\tau}$ must be preserved as in the observed rank vector $\tilde{\bm{\tau}}$, whereas the ranks of the remaining items $a_{j}\in A^{c}$ are left unspecified.
@@
$p(\mathbf{z}_{t}|\tau_{1:t})\propto\mathbb{I}(\mathbf{z}_{t}\in A_{t})\sum_{\mathbf{k}\in\{1,\ldots,K\}^{N}}\mathcal{N}_{N}(\mathbf{z}_{t}|\bm{\tilde{\mu}}_{\mathbf{k}},I_{N})q_{\mathbf{k},t}(A_{1:t-1}),$ (7)
@@
$q_{\mathbf{k},t}(A_{1:t-1})$ $=\sum_{\mathbf{m}\in\{1,\ldots,K\}^{N}}q_{\mathbf{m},t-1}(A_{1:t-2})\int_{\mathbf{C}_{\mathbf{k},t-1}}\mathcal{N}_{N}(\mathbf{z}_{t-1}|\bm{\tilde{\mu}}_{\mathbf{m}},I_{N})\>\mathrm{d}\mathbf{z}_{t-1},$
@@
$p(\mathbf{z}_{t+1}|\tau_{1:t})\propto\sum_{\mathbf{k}\in\{1,\ldots,K\}^{N}}\mathcal{N}_{N}(\mathbf{z}_{t+1}|\bm{\tilde{\mu}}_{\mathbf{k}},I_{N})q_{\mathbf{k},t+1}(A_{1:t}).$ (8)
@@
Corollary 1 shows that both the filtering and smoothing distributions are proportional to finite mixtures. The alternative representation for the filtering distribution and mixture distributions for backward sampling (see the Supplement) can be used to design a more efficient algorithm for posterior inference based on a data augmentation approach coupled with a Metropolis-Hastings step (see Section 3).
@@
$p(\mathbf{z}_{t}|\tau_{1:t})$ $\propto\sum_{\mathbf{k}\in\{1,\ldots,K\}^{N}}w^{F}_{\mathbf{k},t}u^{F}_{\mathbf{k},t}(\mathbf{z}_{t}),$ (9)
@@
$w^{F}_{\mathbf{k},t}$ $=\frac{q_{\mathbf{k},t}(A_{1:t-1})n_{\mathbf{k},t}}{\sum_{\mathbf{m}\in\{1,\ldots,K\}^{N}}q_{\mathbf{m},t}(A_{1:t-1})n_{\mathbf{m},t}}.$
@@
$p(\mathbf{z}_{t}|\tau_{1:T})$ $\propto\sum_{\mathbf{k}\in\{1,\ldots,K\}^{N}}w^{S}_{\mathbf{k},t}u^{S}_{\mathbf{k},t}(\mathbf{z}_{t}),$ (10)
@@
$w^{S}_{\mathbf{k},t}$ $=\frac{q_{\mathbf{k},t}(A_{1:t-1})m_{\mathbf{k},t}}{\sum_{\bm{\ell}\in\{1,\ldots,K\}^{N}}q_{\bm{\ell},t}(A_{1:t-1})m_{\bm{\ell},t}},\qquad\mathbf{k}\in\{1,\ldots,K\}^{N}.$
@@
Furthermore, $p(\mathbf{z}_{1:T}|\tau_{1:T})=p(\mathbf{z}_{T}|\tau_{1:T})\prod_{t=1}^{T-1}p(\mathbf{z}_{t}|\mathbf{z}_{t+1},\tau_{1:(T-1)})$ and backward sampling from $p(\mathbf{z}_{t}|\tau_{1:T},\mathbf{z}_{(t+1):T})$, consists of samples from mixtures of truncated Gaussian distributions. Details are provided in the Supplement.